OpenStudy (anonymous):

Given that sin(x)=1/5 and sin(y)= 5/6, where x and y are acute angles, determine the exact value of sin (X + y)

OpenStudy (mertsj):

\[\sin (x+y)=sinxcosy+cosxsiny\]

OpenStudy (mertsj):

\[\sin (x+y)=(\frac{1}{5})(\frac{\sqrt{11}}{6})+(\frac{2\sqrt{6}}{5})(\frac{5}{6})\]

OpenStudy (mertsj):

\[\sin(x+y)=\frac{\sqrt{11}}{30}+\frac{10\sqrt{6}}{30}=\frac{\sqrt{11}+10\sqrt{6}}{30}\]

OpenStudy (anonymous):

Thank you very much kind sir :)